OpenStudy (anonymous):

simplify 11 cubed root of 6 + 8 fourth root of 6 + 2 cubed root of 6 - 3 fourth root of 6? a) 5 4th root of 6 + 13 3rd root of 6 b) 5 4th root of 12 + 13 3rd root of 12 c) 18√12 d) 18√6

OpenStudy (anonymous):

\[11\sqrt[3]{6}+8\sqrt[4]{6}+2\sqrt[3]{6}-3\sqrt[4]{6}=13\sqrt[3]{6}+5\sqrt[4]{6}\]